# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 3. Harish-Chandra bimodules
$x\xi:=\xi x-\hbar\mathrm{ad}_{\xi}(x),\ x\in X,\xi\in\mathfrak{g}.$
@@
(3.1) $\mathrm{Hom}_{\mathrm{HC}_{\hbar}(\mathrm{GL})}(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes V,\mathrm{U}_{\hbar}(\mathfrak{g})\otimes W)\cong\mathrm{Hom}_{\mathrm{Rep}(G)}(V,\mathrm{U}_{\hbar}(\mathfrak{g})\otimes W),$
@@
The functor $\mathrm{free}$ is monoidal, i.e. for every $V,W\in\mathrm{Rep}(G)$, we have natural isomorphisms
@@
$(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A)\otimes(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A)\rightarrow(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A)\otimes_{\mathrm{U}_{\hbar}(\mathfrak{g})}(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A)\cong\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A\otimes A\rightarrow\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A,$
@@
$a\xi=\xi a-\hbar\mathrm{ad}_{\xi}(a),a\in A,\ \xi\in\mathfrak{g}.$
@@
${}^{R}\iota_{\alpha}^{*}\colon\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}(\mathbf{C}^{N})^{*}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\mathcal{O}(\mathrm{GL}_{N})\cong\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),$
that we denote by the same letter. Similarly, one can extend their left versions (1.2) by identifying $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$ with right-invariant vector fields.
@@
In what follows, we will use the following trick to relate the results of [Mol02] to the setting of the paper. Denote by $\alpha=N+1$. Then there is an algebra map
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1}),\qquad E_{ij}\mapsto E_{ij},\qquad v_{i}\mapsto E_{i\alpha}.$
for all $i,j=1\ldots N$. Likewise, there is an algebra map
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{\bullet}(\mathbf{C}^{N})^{*}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1}),\qquad E_{ij}\mapsto E_{ij},\qquad\phi_{i}\mapsto E_{\alpha i}.$
@@
$\varpi_{N}\colon\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}\xrightarrow{\sim}[\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}(\mathbf{C}^{N})^{*}]^{\mathrm{opp}},\qquad E_{ij}\mapsto E_{ji},\qquad v_{i}\mapsto\phi_{i},$
@@
3.1. Action of the degenerate affine Hecke algebra
@@
(3.2) $\mathbf{Z}_{\mathfrak{g}}\rightarrow\mathrm{End}_{\mathrm{HC}_{\hbar}(G)}(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes V),\qquad z\mapsto(x\otimes v\mapsto zx\otimes v).$
@@
(3.3) $\Omega\colon\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N},\qquad xv_{i}\mapsto\sum_{j}xE_{ij}v_{j}.$
is an endomorphism of Harish-Chandra bimodules. It will be convenient to present it in a matrix form:
(3.4) $(\Omega(v_{1}),\ldots,\Omega(v_{N}))^{T}=E\cdot(v_{1},\ldots,v_{N})^{T},$
@@
Recall the quantum determinant $A(u)$ of (1.11) which we can consider as a polynomial with coefficients in $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})$. Since $\Omega$ commutes with the action of the center, substitution of the variable $u$ with $\Omega$ is a well-defined operation.
@@
By matrix form of (3.4), we need to show that $A(E)=0$. This is [Mol07, Theorem 7.2.1]. ∎
Recall the degenerate affine Hecke algebra $\mathfrak{H}^{\kappa}_{k}$ of ˜2.5. According to [Che98, Example 2.1], there is an action of $\mathfrak{H}^{\hbar}_{k}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$. Recall that there is an isomorphism
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}\cong(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\ldots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}).$
@@
(3.5) $\begin{split}\Omega_{i}:=1\otimes\ldots\otimes\Omega\otimes\ldots\otimes 1\curvearrowright(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N})\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\ldots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}),\\ \sigma_{i}(x\otimes v_{a_{1}}\otimes\ldots\otimes v_{a_{i}}\otimes v_{a_{i+1}}\otimes\ldots\otimes v_{a_{k}}):=x\otimes v_{a_{1}}\otimes\ldots\otimes v_{a_{i+1}}\otimes v_{a_{i}}\otimes\ldots\otimes v_{a_{k}},\end{split}$
where $\Omega$ acts on the $i$-th factor and $\sigma_{i}$ is just a permutation of $(i,i+1)$-th factor on $(\mathbf{C}^{N})^{\otimes k}$ extended to $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$.
Recall the degenerate affine Hecke algebra $\mathfrak{H}^{\hbar}_{k}$ from ˜2.5. The following can be checked directly.
@@
defines an action of $\mathfrak{H}^{\hbar}_{k}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$.
For the following corollary, we use the embedding of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ into $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$ as in ˜3.4.
@@
The action of Proposition˜3.8 induces an action of $\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ such that
$f(\vec{Y})\cdot(v_{i_{1}}\cdot\ldots\cdot v_{i_{k}})=\sum_{J}c_{J}\Omega_{1}^{j_{1}}(v_{i_{1}})\cdot\ldots\cdot\Omega_{k}^{i_{k}}(v_{i_{k}}),\qquad f(\vec{Y})=\sum_{J=(j_{1},\ldots,j_{k})}c_{J}Y_{1}^{j_{1}}\ldots Y_{k}^{j_{k}}.$
@@
Observe that the action is well-defined since $f(\vec{Y})$ is symmetric. The spherical subalgebra $e^{\hbar}_{k}\mathfrak{H}^{\hbar}_{k}e^{\hbar}_{k}$ acts on the invariants $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}]^{S_{k}}$ that, by definition, can be identified with $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$. The explicit formula follows from Proposition˜3.8. ∎
4. Parabolic restriction and difference operators
In this section, we recall the notion of the parabolic restriction for Harish-Chandra bimodules essentially following [KS22], as well as the classical construction of Mickelsson algebras [Zhe90]. In particular, we identify an example of the latter with a negative half of the algebra of difference operators.
4.1. Harish-Chandra bimodules for torus
In this subsection, we explicitly describe the category $\mathrm{HC}_{\hbar}(H_{N})$ and its generic version $\mathrm{HC}_{\hbar}(H_{N})^{\mathrm{gen}}$ that we will use for the parabolic restriction.
Since $H_{N}$ is commutative, the algebra $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})$ can be identified with the functions $\mathcal{O}(\mathfrak{h}_{N}^{*}\times\mathbb{A}^{1})$, where $\mathbb{A}^{1}$ corresponds to $\hbar$. We will mostly use $\rho_{N}$-shifted coordinates (1.4) on $\mathfrak{h}_{N}^{*}\times\mathbb{A}^{1}$
(4.1) $w_{i}:=E_{ii}+N\hbar-i\hbar\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})\subset\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}.$
@@ 4.1. Harish-Chandra bimodules for torus
$f(w):=f(w_{1},\ldots,w_{N},\hbar)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}.$
@@
(4.2) $\mathrm{HC}_{\hbar}(H_{N})\cong\bigoplus_{\lambda\in\Lambda}\mathrm{Mod}_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}[\lambda],$
@@
(4.3) $x\cdot f(w):=f(w-\hbar\lambda)\cdot x,\qquad x\in X,\ f(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N}).$
@@
(4.4) $\mathrm{Mod}_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}[\lambda]\times\mathrm{Mod}_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}[\mu]\rightarrow\mathrm{Mod}_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}[\lambda+\mu],\qquad(X,Y)\mapsto X\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}Y,$
@@
(4.5) $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}:=\mathbf{C}[\hbar][w_{1},\ldots,w_{N}]\left[\frac{1}{w_{i}-w_{j}+k\hbar}\Big{|}\begin{array}[]{cc}i,j=1\ldots N\\ k\in\mathbb{Z}\end{array}\right],$
@@
4.2. Parabolic restriction
In this subsection, we recall the notion of parabolic restriction as in [KS22] and explicitly compute in some concrete examples.

### v4

@@ 4 Harish-Chandra bimodules
$x\xi:=\xi x-\hbar\mathrm{ad}_{\xi}(x),\qquad x\in X,\ \xi\in\mathfrak{g}.$
@@
$\mathrm{Hom}_{\mathrm{HC}_{\hbar}(\mathrm{GL})}(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes V,\mathrm{U}_{\hbar}(\mathfrak{g})\otimes W)\cong\mathrm{Hom}_{\mathrm{Rep}(G)}(V,\mathrm{U}_{\hbar}(\mathfrak{g})\otimes W),$ (4.1)
@@
The functor $\mathrm{free}$ is monoidal, i.e., for every $V,W\in\mathrm{Rep}(G)$, we have natural isomorphisms
@@
$(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A)\otimes(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A)\rightarrow(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A)\otimes_{\mathrm{U}_{\hbar}(\mathfrak{g})}(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A)$
$\qquad{}\cong\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A\otimes A\rightarrow\mathrm{U}_{\hbar}(\mathfrak{g})\otimes A,$
@@
$a\xi=\xi a-\hbar\mathrm{ad}_{\xi}(a),\qquad a\in A,\ \xi\in\mathfrak{g}.$
@@
${}^{R}\iota_{\alpha}^{*}\colon\ \mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\mathcal{O}(\mathrm{GL}_{N})\cong\mathrm{D}_{\hbar}(\mathrm{GL}_{N}),$
that we denote by the same letter. Here $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ is the Rees algebra of differential operators on $\mathrm{GL}_{N}$, see Section 2.3. Similarly, one can extend their left versions (2.2) by identifying $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$ with right-invariant vector fields.
@@
In what follows, we will use the following trick to relate the results of [37] to the setting of the paper. Denote by $\alpha=N+1$. Then there is an algebra map
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1}),\qquad E_{ij}\mapsto E_{ij},\qquad v_{i}\mapsto E_{i\alpha}$
for all $i,j=1,\dots,N$. Likewise, there is an algebra map
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1}),\qquad E_{ij}\mapsto E_{ij},\qquad\phi_{i}\mapsto E_{\alpha i}.$
@@
$\varpi_{N}\colon\ \mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}\xrightarrow{\sim}\bigl[\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr]^{\mathrm{opp}},\qquad E_{ij}\mapsto E_{ji},\qquad v_{i}\mapsto\phi_{i},$
@@
4.1 Action of the degenerate affine Hecke algebra
@@
$\mathbf{Z}_{\mathfrak{g}}\rightarrow\mathrm{End}_{\mathrm{HC}_{\hbar}(G)}(\mathrm{U}_{\hbar}(\mathfrak{g})\otimes V),\qquad z\mapsto(x\otimes v\mapsto zx\otimes v).$
@@
$\Omega\colon\ \mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N},\qquad xv_{i}\mapsto\sum_{j}xE_{ij}v_{j},$ (4.2)
is an endomorphism of Harish-Chandra bimodules. It will be convenient to present it in a matrix form
$(\Omega(v_{1}),\dots,\Omega(v_{N}))^{\mathsf{T}}=E\cdot(v_{1},\dots,v_{N})^{\mathsf{T}},$ (4.3)
@@
Recall the quantum determinant $A(u)$ of (2.8) which we can consider as a polynomial with coefficients in $\mathrm{End}_{\mathrm{HC}_{\hbar}(\mathrm{GL}_{N})}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigr)$. Since $\Omega$ commutes with the action of the center, substitution of the variable $u$ with $\Omega$ is a well-defined operation.
@@
According to matrix form of (4.3), we need to show that $A(E)=0$. This is [38, Theorem 7.2.1]. ∎
Recall the degenerate affine Hecke algebra $\mathfrak{H}^{\kappa}_{k}$ of Definition 3.5. According to [9, Example 2.1], there is an action of $\mathfrak{H}^{\hbar}_{k}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$. Recall that there is an isomorphism
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}\cong\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigr)\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\dots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigr).$
@@
$\Omega_{i}:=1\otimes\dots\otimes\Omega\otimes\dots\otimes 1\curvearrowright\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigr)\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\dots\otimes_{\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})}\bigl(\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}\bigr),$
$\sigma_{i}(x\otimes v_{a_{1}}\otimes\dots\otimes v_{a_{i}}\otimes v_{a_{i+1}}\otimes\dots\otimes v_{a_{k}}):=x\otimes v_{a_{1}}\otimes\dots\otimes v_{a_{i+1}}\otimes v_{a_{i}}\otimes\dots\otimes v_{a_{k}},$ (4.4)
where $\Omega$ acts on the $i$-th factor and $\sigma_{i}$ is just a permutation of $(i,i+1)$-th factor on $\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$ extended to $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$.
Recall the degenerate affine Hecke algebra $\mathfrak{H}^{\hbar}_{k}$ from Definition 3.5. The following can be checked directly.
@@
defines an action of $\mathfrak{H}^{\hbar}_{k}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$.
For the following corollary, we use the embedding of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ into $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$ as in Definition 4.4.
@@
The action of Proposition 4.8 induces an action of $\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]^{S_{k}}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ such that
$f\bigl(\vec{Y}\bigr)\cdot(v_{i_{1}}\cdots v_{i_{k}})=\sum_{J}c_{J}\Omega_{1}^{j_{1}}(v_{i_{1}})\cdots\Omega_{k}^{i_{k}}(v_{i_{k}}),\qquad f\bigl(\vec{Y}\bigr)=\sum_{J=(j_{1},\dots,j_{k})}c_{J}Y_{1}^{j_{1}}\cdots Y_{k}^{j_{k}}.$
@@
Observe that the action is well-defined since $f\bigl(\vec{Y}\bigr)$ is symmetric. The spherical subalgebra ${\rm e}^{\hbar}_{k}\mathfrak{H}^{\hbar}_{k}{\rm e}^{\hbar}_{k}$ acts on the invariants $\bigl[\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}\bigr]^{S_{k}}$ that, by definition, can be identified with $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$. The explicit formula follows from Proposition 4.8. ∎
5 Parabolic restriction and difference operators
In this section, we recall the notion of the parabolic restriction for Harish-Chandra bimodules essentially following [28], as well as the classical construction of Mickelsson algebras [49]. In particular, we identify the one associated to the symmetric algebra $S^{\bullet}\mathbf{C}^{N}$ with a negative half of the algebra of difference operators, see Theorem 5.11. It will be used in Section 7 to define a quantum Toda lattice counterpart of the GKLO homomorphism of Theorem 3.16, one of the key steps for the main result of the paper Theorem 8.10.
5.1 Harish-Chandra bimodules for torus
Let $H_{N}\subset\mathrm{GL}_{N}$ be the subgroup of diagonal matrices as in Section 2.1. In this subsection, we explicitly describe the category $\mathrm{HC}_{\hbar}(H_{N})$ and its generic version $\mathrm{HC}_{\hbar}(H_{N})^{\mathrm{gen}}$ that we will use for the parabolic restriction.
Since $H_{N}$ is commutative, the algebra $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})$ can be identified with the functions $\mathcal{O}\bigl(\mathfrak{h}_{N}^{*}\!\times\!\mathbb{A}^{1}\bigr)$, where $\mathbb{A}^{1}$ corresponds to $\hbar$. We will mostly use $\rho_{N}$-shifted coordinates (2.3) on $\mathfrak{h}_{N}^{*}\times\mathbb{A}^{1}$
$w_{i}:=E_{ii}+N\hbar-i\hbar\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})\subset\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}.$
@@ 5.1 Harish-Chandra bimodules for torus
$f(w):=f(w_{1},\dots,w_{N},\hbar)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}.$
@@
$\mathrm{HC}_{\hbar}(H_{N})\cong\bigoplus_{\lambda\in\Lambda}\mathrm{Mod}_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}[\lambda],$ (5.1)
@@
$x\cdot f(w):=f(w-\hbar\lambda)\cdot x,\qquad x\in X,\ f(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N}).$ (5.2)
@@
$\mathrm{Mod}_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}[\lambda]\times\mathrm{Mod}_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}[\mu]\rightarrow\mathrm{Mod}_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}[\lambda+\mu],\qquad(X,Y)\mapsto X\otimes_{\mathrm{U}_{\hbar}(\mathfrak{h}_{N})}Y,$ (5.3)
@@
$\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}:=\mathbf{C}[\hbar][w_{1},\dots,w_{N}]\biggl[\frac{1}{w_{i}-w_{j}+k\hbar}\>\bigg|\>i,j=1,\dots,N,\,k\in\mathbb{Z}\biggr],$ (5.4)
@@
5.2 Parabolic restriction
In this subsection, we recall the notion of parabolic restriction as in [28] and explicitly compute in some concrete examples.